# LLM-Assisted Static Analysis for Detecting Security Vulnerabilities

### v2

@@ 5.3 RQ2: Previously Unknown Vulnerabilities by IRIS
We applied IRIS with GPT-4 to the latest versions of 30 Java projects. Among the 16 inspected projects where IRIS raised at least one alert, we identified 6 potential vulnerabilities, of which 4 have been reported to the developers and are pending confirmation. These reported vulnerabilities include 3 instances of path injection (CWE-22) and one case of cross-site scripting (CWE-94). To ensure that these vulnerabilities were indeed uncovered due to IRIS’s integration with LLMs, we verified that they were not detectable by CodeQL alone. Detailed findings are presented in the appendix, but we highlight one such vulnerability in Fig. 8. CodeQL was unable to detect this issue due to a missing source specification, while GPT-4 successfully flagged the API endpoint restoreFromCheckpoint as a potential entry point for attack.
@@ 6 Related Work
ML-based approaches for vulnerability detection. Numerous prior techniques incorporate deep learning for detecting vulnerabilities. This includes techniques that use Graph Neural Network (GNN)-based representations of code such as Devign [49], Reveal [3], LineVD [19], and IVDetect [30]; LSTM-based models for representing program slices and data dependencies such as VulDeePecker [32] and SySeVR [31]; and fine-tuning of Transformer-based models such as LineVul [13], DeepDFA [41], and ContraFlow [5]. These approaches focus on method-level detection of vulnerabilities and provide only a binary label classifying a method as vulnerable or not. In contrast, IRIS performs whole-project analysis and provides a distinct code path from a source to a sink and can be tailored for detecting different CWEs. More recently, multiple studies demonstrated that LLMs are not effective at detecting vulnerabilities in real-world code [42, 10, 25]. While these studies only focused on method-level vulnerability detection, it reinforces our motivation that detecting vulnerabilities requires whole-project reasoning, which LLMs currently cannot do alone.
Static analysis tools. Apart from CodeQL [1], other static analysis tools like CppCheck [8], Semgrep [37], FlawFinder [12], Infer [11], and CodeChecker [7] also include analyses for vulnerability detection. But, these tools are not as feature-rich and effective as CodeQL [29, 33]. Recently, proprietary tools such as Snyk [39] and SonarQube [40] are also gaining in popularity. However, like CodeQL, these tools share the same fundamental limitations of missing specifications and false positives, which IRIS improves upon. Potentially, our techniques stand to benefit all such tools.
LLM-based approaches for software engineering. Researchers are increasingly combining LLMs with program reasoning tools for challenging tasks such as fuzzing [27, 46], program repair [45, 22, 44], and fault localization [47]. While we are on a similar direction as [28], to our knowledge, our work is among the first to combine LLMs with static analysis to detect application level security vulnerabilities via whole-project analysis. Recently, LLM-based agents such as AutoCodeRover [48] and SWE-Agent [43] are also pushing the boundaries on whole-project repair. Hence, in future, we plan to explore a richer combination of tools in IRIS to further improve the performance of vulnerability detection.
@@ A.1 Selecting Candidate Specifications
We note that this is not the entire story as there might be other kinds of sources and sinks. Other types of source candidates include the formal parameter of protected but overridden internal functions (the req parameter in protected HTTPServeletResponse doGet(HTTPServeletRequest req)), arguments to an impure external function (the buffer argument to void read(byte[] buffer, int size)), etc. Sink candidates include the return value of public facing functions, thrown exceptions, and even static methods without any parameter (System.exit()). Due to the complexity, we do not tackle such kind of sources of sinks in this work. However, we plan to explore further in future works.
@@ A.2 LLM Prompts for Specification Inference
Listing 1: LLM prompt for labelling external APIs as sources or sinks
@@
A.3 CodeQL Queries for Static Analysis
@@
Listing 3: QL Script for Detecting Vulnerabilities for Path Traversal (CWE 22)
Listing 3 presents our CodeQL query for Path Traversal vulnerability (CWE 22). In Listing 3, Lines 10-29 describe a taint analysis configuration that describes which nodes in the data flow graph should be considered as sources or sinks. Here, Line 12 specifies our custom predicate isLLMDetectedSource that checks whether the method called is taint source based on our specifications. Similarly, our predicates isLLMDetectedSink checks whether the node is a taint sink based on our specifications. Line 16 checks if a method call or method argument node is a taint sink based on our specifications. We generate the source and sink specifications as predicates in QL file as shown in Listings 4 and 5 respectively. Given a taint configuration and the source and sink specifications, CodeQL can automatically perform taint analysis on a given project.
We use template to convert LLM inferred specifications into CodeQL queries. There are three kinds of queries:
@@
    the return value of an external function as a source;
@@
Listing 4: QL Predicates for Source Specifications
@@
Listing 5: QL Predicates for Sink Specification
A.4 Visualization of Metrics
@@
$\begin{array}[]{rlrl}\footnotesize\textit{VulPaths}(P)=&\{\textit{Path}\in\textit{Paths}^{P}~{}|~{}\textit{Path}~{}\cap~{}\mathbf{V}_{\text{vul}}^{P}\neq\emptyset\}\end{array}$
@@ B.1 Details of Dataset Extraction Process
Table 4: Vulnerability Dataset Collection Statistics
@@
The closest dataset to ours, in terms of features, is the Java dataset curated by Li et al. [29], containing 165 CVEs. While we initially considered using their dataset for our work, we found several issues. First, their dataset does not come with build scripts, which makes it hard to automatically run each project with CodeQL. Second, their dataset only has few CVEs for all but one CWE, which makes it difficult to thoroughly analyze a tool for different vulnerability classes. Finally, they do not provide any automated scripts to curate more CVEs. Hence, we curated our own dataset and our framework also allows to easily extend to more vulnerability classes.
@@ B.2 Comparison of our CWE-Bench-Java with Existing Vulnerability Datasets
    CVE Metadata: whether CVE Metadata is contained in the dataset;
@@
    Compilable: whether the dataset ensures that the projects are end-to-end and automatically compilable;
@@
Table 5: Comparison of CWE-Bench-Java with existing vulnerability datasets
Dataset Languages CVE Metadata Real-World Fix Locations Compilable Vetted
@@
Li et al. [29] Java
SVEN [17] C++
Our Dataset Java
@@ C.1 Experimental Settings
We also select instruction-tuned versions of six state-of-the-art open-source LLMs via huggingface API: Llama 3 8B and 70B, DeepSeekCoder 7B and 33B, Mistral 7B, and Gemma 7B. To run the open-source LLMs we use two groups of machines: a 2.50GHz Intel Xeon machine, with 40 CPUs, four GeForce RTX 2080 Ti GPUs, and 750GB RAM, and another 3.00GHz Intel Xeon machine with 48 CPUs, 8 A100s, and 1.5T RAM.
@@ C.4 Details of Selected LLMs
We include the versions of selected LLMs in Table 6.
Table 6: Selected LLM Versions
@@
C.5 Previously Unknown Bugs Found by IRIS
We have applied IRIS with GPT-4 to 30 open-source Java projects in order to find previously unknown vulnerabilities. The detailed statistics, at the time of this paper’s submission, is listed in Table 7. We inspect vulnerabilities in detail, and will only file when we think that there is a vulnerability with high severity and an easily accessible attack surface. 3 out of the 4 filed vulnerabilities are filed on HackerOne to abide to the rules of the individual open-source project, which prohibit us to disclose the details before the vulnerability is fixed. The remaining one is found in the Java project alluxio 2.9.4, which we disclose and discuss in detail in Fig. 8. For the already filed vulnerabilities, we are committed to work till the end to make sure that they are fixed in the open-source project. The paper will be updated when they are fixed by the developer.
Table 7: Statistics related our effort to identify previously unknown vulnerabilities with IRIS + GPT-4, to the date of this paper’s submission.
#Projects 30
#Projects where IRIS reports $>0$ alarm 16
Average #alarms per investigated project 5.78
#Projects with potential vul 4
#Vuls with potential attack surface 6
#Vuls filed for developer confirmation 4
#Vuls of CWE-22 3
#Vuls of CWE-78 0
#Vuls of CWE-79 0
#Vuls of CWE-94 1
#Developer Response 0
#Developer Confirmation 0
#Fixed Vulnerability 0
Our experience with finding unknown vulnerabilities using IRIS is pleasant, especially given that we have built a custom web-based interface for us to inspect and triage the vulnerabilities. A screenshot of the user-interface is illustrated in Fig. 11. We may quickly scan over the vulnerabilities by looking at their source, sink, and LLM explanation of whether it is a bug. Additionally, all the dataflows and nodes have hyperlinks to Github and the source code, making it easy to navigate through the codebase. Overall, it does not take too much time to dive deep into each vulnerability. Anecdotally, the first two real CWE-22 vulnerabilities are found among the first two projects that we inspected.
Figure 11: Our web-based user-interface for inspecting and triaging vulnerabilities. The screenshot shows the detected known vulnerability CVE-2021-41269 which we also detail in our motivating example section. For detecting unknown vulnerability, the information about CWE and golden labels will not be available to the users.
C.6 Statistics of Unique and Recurring Specifications
@@
C.7 Statistics of Inferred Taint Specifications
@@ Appendix D Analysis Runtime
We include the full table containing statistics to provide more details about projects and our analysis (Table LABEL:tab:longtable). For each project, we present its corresponding CWE ID, the lines-of-code (SLOC), the time it takes to run the full analysis, the number candidate APIs and the number of labeled source and sinks by Llama 3 8B. We also color code cells of interest: For SLOC, we mark a cell as red if $>$1M; yellow if $>$100k. For Time, we mark a cell as red if $\geq$1h; yellow if $\geq$5m. For the number of candidates, we mark a cell as red if $>$10k. Lastly for sources and sinks, we mark a cell as red if the number is larger than 200.
Table 11: Details of analysis runtime, candidates, and inferred sources and sinks for all projects (Llama 3 8B)

### v3

@@ 5.3 RQ2: Previously Unknown Vulnerabilities by IRIS
We applied IRIS with GPT-4 to the latest versions of 30 Java projects. Among the 16 inspected projects where IRIS raised at least one alert, we identified 4 vulnerabilities, including 3 instances of path injection (CWE-22) and one case of code-injection (CWE-94). To ensure that these vulnerabilities were indeed uncovered due to IRIS’s integration with LLMs, we verified that CodeQL alone did not detect them. We highlight one such vulnerability in Fig. 8. CodeQL was unable to detect this issue due to a missing source specification, while GPT-4 successfully flagged the API endpoint restoreFromCheckpoint as a potential entry point for attack.
@@ 6 Related Work
Learning-based approaches for vulnerability detection. Numerous prior techniques incorporate deep learning for detecting vulnerabilities. This includes Graph Neural Network based models such as Zhou et al. (2019); Chakraborty et al. (2020); Dinella et al. (2020); Hin et al. (2022); Li et al. (2021a); LSTM-based models such as Li et al. (2020; 2021b); and fine-tuning of Transformer-based models such as Fu & Tantithamthavorn (2022); Steenhoek et al. (2023); Cheng et al. (2022). These approaches focus on method-level detection of vulnerabilities and provide only a binary label classifying a method as vulnerable or not. In contrast, IRIS performs whole-project analysis and provides a distinct code path from a source to a sink and can be tailored for detecting different CWEs. More recently, multiple studies demonstrated that LLMs are not effective at detecting vulnerabilities in real-world code (Steenhoek et al., 2024; Ding et al., 2024; Khare et al., 2023). While these studies only focused on method-level vulnerability detection, it reinforces our motivation that detecting vulnerabilities requires whole-project reasoning, which LLMs currently cannot do alone.
Static analysis tools. Apart from CodeQL (Avgustinov et al., 2016), other static analysis tools (CPPCheck, ; Semgrep, 2023; FlawFinder, ; FB Infer, ; Code Checker, ) also include analyses for vulnerability detection. More general query engines (Scholz et al., 2016; Li et al., 2023b) have also been applied to find program bugs. But these tools are not as feature-rich and effective as CodeQL (Li et al., 2023a; Lipp et al., 2022). Recently, proprietary tools such as Snyk (Snyk.io, ) and SonarQube (SonarQube, ) are also gaining in popularity, although sharing the same fundamental limitations of missing specifications and false positives, which IRIS improves upon. We envision our technique to benefit all such tools. Works such as MERLIN (Livshits et al., 2009), Seldon (Chibotaru et al., 2019) and InspectJS (Dutta et al., 2022) tackle the problem of specification inference through probabilistic modeling. Specifically, like IRIS, InspectJS also augments CodeQL with specifications inferred using machine learning. However, InspectJS relies on the quality of seed specifications and requires expensive analysis of each third-party library, which IRIS does not—making it more scalable. Future work could explore incorporating probability estimates for specifications.
LLM-based approaches for software engineering. Researchers are increasingly combining LLMs with program reasoning tools for challenging tasks such as fuzzing (Lemieux et al., 2023; Xia et al., 2024), program repair (Xia et al., 2023; Joshi et al., 2023; Xia & Zhang, 2022), and fault localization (Yang et al., 2023). While we are on a similar direction as (Li et al., 2024; Wang et al., 2024), to our knowledge, our work is among the first to combine LLMs with static analysis to detect application-level security vulnerabilities via whole-project analysis.
@@ 7 Conclusion and Limitations
Acknowledgments
We thank the anonymous reviewers for their valuable feedback and suggestions that helped improve this work. We also thank Claire Wang for helping with open-sourcing the project. This research was supported by NSF award CCF 2313010.
@@ A.1 Selecting Candidate Specifications
We note that this is not the entire story as there might be other kinds of sources and sinks. Other types of source candidates include the formal parameter of protected but overridden internal functions (the req parameter in protected HTTPServeletResponse doGet(HTTPServeletRequest req)), arguments to an impure external function (the buffer argument to void read(byte[] buffer, int size)), etc. Sink candidates include the return value of public facing functions, thrown exceptions, and even static methods without any parameter (System.exit()). Due to the complexity, we do not tackle such kind of sources of sinks in this work. However, we plan to explore further in future work.
@@ A.2 LLM Prompts for Specification Inference
Listing 1: LLM prompt for labelling external APIs as sources or sinks.
@@
A.3 Design Decisions of LLM Prompts
Here are some reasoning behind the design decisions for the LLM prompts:
  1. 1.
    Few-shot for labeling external API: Since the LLM is tasked with labeling source, sink, and taint-propagator APIs, we provide one example for each category, along with a negative example of an API that does not fall into any of these categories. We typically select examples from the Java standard libraries because they are widely used and their labels are readily available.
  2. 2.
    Zero-shot for labeling internal API: As labels for internal APIs are not available, we rely on the zero-shot capabilities of the language model. To mitigate potential performance loss, we include additional information, such as documentation associated with important internal APIs.
  3. 3.
    ±$\pm 5$ lines surrounding the source and sink location during contextual analysis: We chose ±$\pm 5$ lines as a balanced approach to provide sufficient context while managing performance and cost. While technically possible to use a larger window, we observed that excessive context can overwhelm the language model, leading to reduced accuracy. Additionally, a larger context increases computational costs significantly, particularly given the large number of candidate APIs and paths that must be queried.
  4. 4.
    Selecting subset of nodes in the alarm path during contextual analysis: We use a hyperparameter $S$ to control the number of intermediate steps included in the prompt. For paths with more than $S$ intermediate steps, we divide the path into $S$ equal segments and select one step from each. This selection prioritizes function calls, as they may indicate sanitizations. If no function call is present, a node is randomly selected from the segment. In our experiments, we observe that setting $S$ to 10 provides a good balance between the cost and accuracy so that the prompt contains enough context and would not be too long.
A.4 CodeQL Queries for Static Analysis
@@
Listing 3: CodeQL script for detecting vulnerabilities for Path-Traversal (CWE 22).
Listing 3 presents our CodeQL query for Path-Traversal vulnerabilities (CWE 22). Lines 10-29 describe a taint analysis configuration that describes which nodes in the data flow graph should be considered as sources or sinks. Here, Line 12 specifies our custom predicate isLLMDetectedSource that checks whether the method called is taint source based on our specifications. Similarly, our predicates isLLMDetectedSink checks whether the node is a taint sink based on our specifications. Line 16 checks if a method call or method argument node is a taint sink based on our specifications. We generate the source and sink specifications as predicates in QL file as shown in Listings 4 and 5 respectively. Given a taint configuration and the source and sink specifications, CodeQL can automatically perform taint analysis on a given project.
We use templates to convert LLM inferred specifications into CodeQL queries. There are three kinds of queries:
@@
    the return value of an external function as a source; and
@@
Listing 4: CodeQL predicate for source specifications.
@@
Listing 5: CodeQL predicate for sink specifications.
A.5 Visualization of Metrics
@@
$\begin{array}[]{rlrl}\textit{VulPaths}(P)=&\{\textit{Path}\in\textit{Paths}^{P}~{}|~{}\textit{Path}~{}\cap~{}\mathbf{V}_{\text{vul}}^{P}\neq\emptyset\}\end{array}$
@@ B.1 Details of Dataset Extraction Process
Table 4: Vulnerability dataset collection statistics.
@@
The closest dataset to ours, in terms of features, is the Java dataset curated by Li et al. (2023a), containing 165 CVEs. While we initially considered using their dataset for our work, we found several issues. First, their dataset does not come with build scripts, which makes it hard to automatically run each project with CodeQL. Second, their dataset only has few CVEs for all but one CWE, which makes it difficult to thoroughly analyze a tool for different vulnerability classes. Finally, they do not provide any automated scripts to curate more CVEs. Hence, we curated our own dataset and our framework also allows to easily extend to more vulnerability classes.
@@ B.2 Comparison of our CWE-Bench-Java with Existing Vulnerability Datasets
    CVE Info: whether CVE Metadata is contained in the dataset;
@@
    Compilable: whether the dataset ensures that the projects are end-to-end and automatically compilable; and
@@
Table 5: Comparison of CWE-Bench-Java with existing vulnerability datasets.
Dataset Languages CVE Info Real-World Fix Locations Compilable Vetted
@@
Li et al. (2023a) Java
SVEN (He & Vechev, 2023) C++
CWE-Bench-Java (Ours) Java
@@ C.1 Experimental Settings
We also select instruction-tuned versions of six state-of-the-art open-source LLMs via huggingface API: Llama 3 8B and 70B, DeepSeekCoder 7B, QWEN-2.5-Coder 32B, and Gemma-2 27B. To run the open-source LLMs we use two groups of machines: a 2.50GHz Intel Xeon machine, with 40 CPUs, four GeForce RTX 2080 Ti GPUs, and 750GB RAM, and another 3.00GHz Intel Xeon machine with 48 CPUs, 8 A100s, and 1.5T RAM.
@@ C.4 Details of Selected LLMs
We include the versions of the 7 selected LLMs in Table 6.
Table 6: Versions and model IDs of the selected LLMs in our evaluation.
@@
Qwen-2.5-Coder 32B Instruct Qwen/Qwen2.5-Coder-32B-Instruct
Gemma-2 27B google/gemma-2-27b-it
C.5 Statistics of Unique and Recurring Specifications
@@
C.6 Statistics of Inferred Taint Specifications
@@
C.7 Error Analysis: Cause of Undetected Vulnerabilities
In general, we find the following three main causes of undetected vulnerabilities by employing IRIS:
  1. 1.
    Vulnerability cannot be modeled by simple taint dataflows: for instance, the vulnerability CVE-2020-11977 (CWE-94) is manifested by an unexpected exit(1) call, with no direct taint dataflow going into it. In fact, the taint dataflow goes into the condition of an “if” statement surrounding the exit. In this case, the model of sink API cannot capture the vulnerability.
  2. 2.
    Missing dataflow edge due to side-effects: for instance a taint information is written to a temporary file, and is later read from the same file, subsequently causing a vulnerability. However, the dataflow edge is carried through side-effects, which is not captured by CodeQL.
  3. 3.
    Missing dataflow edge due to unspecified library usage: The vulnerability can only be manifested through a concrete usage of the library; but within the library itself there is no possible dataflow to connect the source and the sink.
Overall, we view the above as general limitations for the static analysis. In terms of LLM induced false negatives, here are the two main failure modes:
  1. 1.
    Missing taint-propagator labels from LLM: this would cause missing dataflow edges stopping source to flow to sink.
  2. 2.
    Missing source or sink specifications from LLM: if there is no relevant source/sink specification then the static analysis tool would not have the anchor for analysis.
@@ Appendix D Analysis Runtime
We include the full table containing statistics to provide more details about projects and our analysis (Table LABEL:tab:longtable). For each project, we present its corresponding CWE ID, the lines-of-code (SLOC), the time it takes to run the full analysis, the number candidate APIs and the number of labeled source and sinks by Llama 3 8B. We also color code cells of interest: For SLOC, we mark a cell as red if $>$1M; yellow if $>$100k. For Time, we mark a cell as red if $\geq$1h; yellow if $\geq$5m. For the number of candidates, we mark a cell as red if $>$10k. Lastly for the numbers of sources and sinks, we mark a cell as red if the number is larger than 200.
Table 10: Details of analysis runtime, candidates, and inferred sources and sinks for all projects (Llama 3 8B).